### v2

# DEFAME: Dynamic Evidence-based FAct-checking with Multimodal Experts

Abstract
The proliferation of disinformation demands reliable and scalable fact-checking solutions. We present Dynamic Evidence-based FAct-checking with Multimodal Experts (DEFAME), a modular, zero-shot MLLM pipeline for open-domain, text-image claim verification. DEFAME operates in a six-stage process, dynamically selecting the tools and search depth to extract and evaluate textual and visual evidence. Unlike prior approaches that are text-only, lack explainability, or rely solely on parametric knowledge, DEFAME performs end-to-end verification, accounting for images in claims and evidence while generating structured, multimodal reports. Evaluation on the popular benchmarks VERITE, AVeriTeC, and MOCHEG shows that DEFAME surpasses all previous methods, establishing itself as the new state-of-the-art fact-checking system for uni- and multimodal fact-checking. Moreover, we introduce a new benchmark, ClaimReview2024+, featuring claims after the knowledge cutoff of GPT-4o, avoiding data leakage. Here, DEFAME drastically outperforms the GPT-4o Chain-of-Thought baseline, demonstrating temporal generalizability and the potential for real-time fact-checking [^2] We released the code and benchmark dataset publicly at:
https://github.com/multimodal-ai-lab/DEFAME
Machine Learning, ICML
1 Introduction
In recent years, misinformation has been growing in scale and quality (Chen & Shu, 2024) beyond human capacity to fact-check. “Fake news” has evolved from a lighthearted term into a serious global threat (World Economic Forum, 2024). Driven by higher engagement rates on social media and the increase in AI use, misinformation spreads faster, reaches a broader audience, and causes greater harm  (Li & Xie, 2020; Zannettou et al., ; Wang et al., ; Chen & Shu, 2024). Humans perceive multimodal information as more credible (Newman et al., 2012; Hameleers et al., 2020), often interpreting visuals as “evidence” (Greifeneder et al., 2021), making it particularly persuasive. Approximately $80\%$ of the claims checked by professional fact-checkers are multimodal (Dufour et al., 2024)—signifying a strong priority for checking multimodal content.
Figure 1: DEFAME in a nutshell: It fact-checks multimodal claims using multimodal evidence and returns a detailed, human-friendly report document.
Unfortunately, Automated Fact-Checking (AFC) systems are mostly text-only (Dmonte et al., 2024; Vykopal et al., 2024). Only a few works venture into visual claim verification (Khaliq et al., 2024; Shao et al., 2023), but none can handle both multimodal claims and multimodal evidence at once. Most multimodal claim verification systems cannot even retrieve the evidence needed to verify a claim (Fu et al., 2024; Vo-Hoang et al., 2024; Tang et al., 2024),. The majority of AFC works focuses on only a specific aspect of AFC such as evidence retrieval (Cheung & Lam, 2023), evidence summarization (Chen et al., 2024), or evidence ranking (Tahmasebi et al., 2024). This specialization has created a scattered landscape where individual approaches address only isolated aspects of a complex problem. AFC systems that target the overall task of fact-checking typically are text-only (Zhao et al., 2024; Li et al., 2024c), lack performance (Chen et al., 2024; Cao et al., 2023; Papadopoulos et al., 2024a), do not involve a flexible planning module (Khaliq et al., 2024; Tahmasebi et al., 2024), or are not explainable (Shao et al., 2023; Xu et al., 2024).
Therefore, we introduce DEFAME, a straightforward end-to-end AFC framework, unifying the advancements in the field of AFC into one single system. As the first of its kind, it is able to natively process multimodal claims and evidence, the latter of which it retrieves dynamically as needed. DEFAME is designed with transparency in mind, imitating a human fact-checking process and returning a detailed fact-check report to the user (cf. Figure 1).
Our contributions can be summarized as follows: First, we present DEFAME, the first multimodal AFC system that can handle multimodal claims as well as retrieve and process multimodal evidence. DEFAME brings together the advancements of previous work by unifying their strengths into one single, yet straightforward framework.
Second, we empirically demonstrate DEFAME’s effectiveness by establishing new state-of-the-art results on three diverse and widely used benchmarks. On AVeriTeC (Schlichtkrull et al., 2024b), we improve accuracy from $65.6\%$ to $74.0\%$; on MOCHEG (Yao et al., 2023), we achieve a $+10.9\%$ improvement in micro F1 score, and on VERITE (Papadopoulos et al., 2024b), we enhance True/False accuracy by $+26.5\%$.
Third, we contribute a new benchmark, ClaimReview2024+, with claims that occurred after the knowledge cutoff of GPT-4o to mitigate the effects of data leakage and show that DEFAME is strongly superior to GPT-4o on these “unseen” statements.
Fourth, we demonstrate that the fact-checking reports generated by DEFAME are well-received by human evaluators, preferring the reports over GPT-4o’s outputs.
Finally, we made our code and benchmark publicly available [^3] https://github.com/multimodal-ai-lab/DEFAME.
2 Related Work
Method Multimodal claims Multimodal evidence Evidence retrieval Multi-hop Planning Tools Reasoning Explainable Open- domain Backbone
RAGAR (Khaliq et al., 2024) - - - - MLLM
MMOOC-Checker (Xu et al., 2024) - - - - - - - VLT
Vo-Hoang et al. (Vo-Hoang et al., 2024) - - - - - - - LLM & VLT
Zeng et al. (Zeng et al., 2024a) - - - - - - - MLLM
CHASMA (Papadopoulos et al., 2024b) - - - - - - - CLIP & VLT
AITR (Papadopoulos et al., 2024a) - - - - - - - VLT
HAMMER (Shao et al., 2023) - - - - - - - - BERT & ViT
MultiMD (Fu et al., 2024) - - - - - - - - VLT
LVLM4FV (Tahmasebi et al., 2024) - - - - (M)LLM
MOCHEG (Yao et al., 2023) - - - - - VLT
MetaSum (Chen et al., 2024) - - - - - VLT & LLM
M³D (Tang et al., 2024) - - - - - - - - VLT & GCN
ChartBERT (Akhtar et al., 2023a) - - - - - - - - BERT
PACAR (Zhao et al., 2024) - - LLM
HiSS (Zhang & Gao, 2023) - - - - LLM
ProgramFC (Pan et al., 2023) - - LLM
Self-Checker (Li et al., 2024c) - - - - LLM
FactLLaMA (Cheung & Lam, 2023) - - - - - - - LLM
CFR (Sriram et al., 2024) - - - - - - - BERT & LLM
GPT-CoT (Cao et al., 2023) - - - - - - LLM
DEFAME (Ours) MLLM
Table 1: Overview of the most relevant and published AFC systems. We consider a method “Explainable” if it offers substantial (✓) or some (◑) human-readable information explaining the decision. Methods with “VLT” backbones employ a model from the Vision Language Transformer family.
Automating the task of fact-checking is a difficult problem, far from being solved (Akhtar et al., 2023b; Dmonte et al., 2024; Yao et al., 2023; Schlichtkrull et al., 2024a). Due to its complexity, Akhtar et al. (2023b) subdivide AFC into three stages: (1) claim detection & extraction, (2) evidence retrieval, and (3) verdict prediction. Most approaches narrow down their scope, either by focusing on a sub-task like summarization (Chen et al., 2024), justification generation  (Atanasova et al., 2020), evidence retrieval  (Samarinas et al., 2021), deepfake detection (Jia et al., 2024), out-of-context detection (Xu et al., 2024; Vo-Hoang et al., 2024), or by addressing only a specific domain, e.g., charts (Akhtar et al., 2023a, 2024), social media (Wang et al., 2018), politics (Khaliq et al., 2024), or news (Xu et al., 2024; Shao et al., 2023). Others investigate the incorporation of new architectures like a knowledge graph (Cao et al., 2024) or address the problem of evidence ambiguity (Glockner et al., 2024). In contrast to all these methods, DEFAME combines the fragmented landscape of AFC work into one end-to-end solution—not restricted to only one modality, domain, or sub-task. Table 1 compares DEFAME to prior work.
Text-only fact-checking
The vast majority of proposed AFC systems is purely text-based (Hassan et al., 2017; Thorne et al., 2018; Schlichtkrull et al., 2024b; Dmonte et al., 2024; Vykopal et al., 2024; Yang & Rocha, 2024; Zhao et al., 2024; Wang et al., 2024b; Li et al., 2024b; Pan et al., 2023; Schlichtkrull et al., 2024a; Cheung & Lam, 2023). The most popular benchmarks to evaluate these methods are LIAR (Wang, 2017), FEVER (Thorne et al., 2018), and AVeriTeC (Schlichtkrull et al., 2024b), the latter of which mitigates important weaknesses of the previous ones, including reliance on artificial claims and evidence insufficiency. For this reason, we use AVeriTeC in our evaluation. Most recently, Factcheck-Bench (Wang et al., 2024a) was introduced to evaluate the factuality of entire LLM responses.
Multimodal fact-checking
Popular multimodal AFC benchmarks include works from Xu et al. (2024); Nielsen & McConville (2022); Zlatkova et al. (2019); Nakamura et al. (2020); Aneja et al. (2021); Yao et al. (2023); Jaiswal et al. (2017); Sabir et al. (2018); Müller-Budack et al. (2020); Luo et al. (2021a). Notably, MOCHEG (Yao et al., 2023) builds on real-world claims, requiring multimodal evidence retrieval, additionally incorporating the task of justification generation. Given these features, we use MOCHEG in our evaluation. A strong emphasis of multimodal AFC has been on Out-Of-Context (OOC) image detection, sometimes referred to as “cheapfake” detection. Popular evaluation benchmarks include NewsCLIPpings (Luo et al., 2021b) and, more recently, VERITE (Papadopoulos et al., 2024b). In our experiments, we use VERITE because it improves over previous benchmarks by reducing unimodal bias and incorporating real-world samples.
Some multimodal AFC solutions utilize multimodal fusion (Papadopoulos et al., 2023, 2024a; Shao et al., 2023; Zeng et al., 2024b), leverage inter-, and cross-modal consistency (Abdelnabi et al., 2022; Fu et al., 2024), or apply conventional machine learning models (Papadopoulos et al., 2024a; Wang et al., 2018). A strong disadvantage of these systems is the inability to produce human-understandable explanations of the predictions. Furthermore, in stark contrast to DEFAME, they often rely on superficial pattern matching or lexical/visual similarity, ignoring factuality and logic, raising questions about their robustness and actuality.
(M)LLM-based fact-checking
With the rise of (Multimodal) Large Language Models or (M)LLMs, the AFC community increasingly explored prompt-driven solutions (Yang & Rocha, 2024; Khaliq et al., 2024; Tahmasebi et al., 2024; Chen et al., 2024; Pan et al., 2023; Cheung & Lam, 2023). One of DEFAME’s closest relatives, RAGAR (Khaliq et al., 2024), processes textual and visual claims but retrieves only textual evidence. Furthermore, DEFAME directly incorporates the claim image in its context while RAGAR converts it into a textual description, discarding critical visual information.
While there have been efforts to improve the performance of MLLM-based approaches through training on synthetic data  (Zeng et al., 2024a), most still rely on the parametric knowledge of the MLLMs (Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e; Geng et al., 2024), foregoing external evidence retrieval. This approach has three major drawbacks: (1) MLLM knowledge is static and fails on recent claims, as shown in this work; (2) predictions lack links to verifiable sources, reducing transparency; and (3) reliance on parametric knowledge increases hallucination risks, making such methods less reliable than Retrieval-Augmented Generation (RAG)-based approaches. In contrast, DEFAME uses internal knowledge only for commonsense reasoning, retrieving all evidence dynamically through external tools. Unlike some prior work (Yang & Rocha, 2024; Tang et al., 2024), it does not rely on benchmark-provided gold evidence, reinforcing its adaptability to real-world misinformation.
3 DEFAME Approach
Large Language Model (LLM) agents have become a powerful solution for commonsense reasoning, summarization, basic planning, and tool use (Davis, 2023; Li et al., 2024d; Zeng et al., 2023; Surís et al., 2023). DEFAME (see Figure 2) comprises a Multimodal LLM (MLLM), a suite of multimodal tools, and a structured fact-check report. Our proposed framework effectively operates as a dynamic, multi-step RAG system (Lewis et al., 2020), inspired by established fact-checking workflows (Moreno Gil et al., 2021). Each call to the MLLM includes the current state of the fact-checking report as contextual input, along with a task-specific description. This approach emulates a form of context awareness, guiding the MLLM to focus on pertinent information at each stage of the fact-checking process and allowing for more intricate, multi-hop reasoning and evidence retrieval. Nonetheless, LLM’s limitations like hallucinations, knowledge cutoff, and stochastic outputs (Maynez et al., 2020; Li & Flanigan, 2024) necessitate careful management. Accordingly, we decompose the fact-checking process into six manageable stages, five of which are subject to MLLM prompting. The procedure mimics human fact-checkers and is described in detail in the following.
Figure 2: Overview of DEFAME’s dynamic six-stage pipeline involving three main components: an MLLM, a fact-checking report, and external tools.
Stage 1: Plan Actions
Upon receiving a claim, the MLLM is prompted to suggest a targeted action sequence to retrieve missing information. Since the action space is infinite for some tools (e.g., web search allows arbitrary queries), the “planner” aims to minimize actions and cost. To prevent redundancy, DEFAME tracks previously executed actions and adapts if it encounters a “dead end.” In-context learning guides the module in deciding which of the specialized tools to invoke: Web Search, Image Search, Reverse Image Search (RIS), or Geolocation. While RIS and Geolocation handle image inputs, Web Search and Image Search operate on dynamically generated text queries.
Stage 2: Execute Actions
Given a set of actions, DEFAME invokes the corresponding tool:
  1. 1.
    Web Search: Given a textual search query, it leverages Google Search via the Serper API [^4] https://serper.dev to provide the top $3$ relevant web pages matching the query. This gives DEFAME foundational access to the open web, enabling it to retrieve current or very domain-specific evidence that is not encoded in the MLLM’s parameters.
  2. 2.
    Image Search: It applies Google Image Search to return up to $3$ URLs of web pages containing images, diagrams, and infographics that match a given textual caption. These retrieved visuals can serve as evidence in the report and/or as inputs for DEFAME’s Geolocation and RIS tools.
  3. 3.
    Reverse Image Search (RIS): For a given image, it retrieves up to 3 URLs of web pages containing the same image, using the Google Vision API [^5] https://cloud.google.com/vision/. It serves as a real-time image knowledge base, enabling DEFAME to contextualize a given image with information such as image sources, authors, prior uses, and approximate dates–—that a static trained MLLM, with fixed parametric knowledge, cannot offer.
  4. 4.
    Geolocation integrates GeoCLIP (Cepeda et al., 2023)—a specialized geolocation tool designed to estimate the most probable countries from which an image could originate. MLLMs lack the ability for this AFC-critical task (Roberts et al., 2024).
To prevent temporal leakage, all web-based tools restrict search results to sources published before the claim’s release date (if known). Additionally, we exclude major fact-checking websites and any sites that disallow automated bot access. A list of excluded domains can be found in Appendix A. For each retrieved URL, we scrape the corresponding page using Firecrawl [^6] https://github.com/mendableai/firecrawl. Unlike previous work, we extend the scraper to identify and download any referenced image, ensuring a complete context for the fact-check.
Stage 3: Summarize Results
At this stage, the gathered evidence is integrated into the fact-checking report, which guides the MLLM’s reasoning through in-context learning. The model generates an abstractive summary of key findings for each tool output, ensuring brevity and alignment with the existing report. Relevant images are retrieved and incorporated, while irrelevant results are filtered by instructing the MLLM to return NONE if they do not contribute meaningfully to the verification process.
Stage 4: Develop the Fact-Check
Corresponding to stage 4 in Moreno Gil et al. (2021), DEFAME brings claim and summarized evidence together. It directs the MLLM to discuss the claim’s veracity step-by-step based on the evidence, flagging any gaps as “incomplete” if the information is missing. This stage offers room for intricate reasoning, deducing new insights through natural language inference.
Stage 5: Predict a Verdict
At this stage, DEFAME assesses the claim’s veracity, classifying it into one of the benchmark-specific categories. The MLLM is prompted to summarize key findings from the report and select a verdict. If there is insufficient information, it may return [Uncaptioned image]NEI. When [Uncaptioned image]NEI is returned, the fact-checking process returns to Stage 1 to gather further evidence. This allows DEFAME to dynamically dive deeper into unresolved questions, allowing for a more thorough exploration of relevant sub-branches. DEFAME moves to the final stage if the end of the third iteration or a definitive verdict is reached, mirroring the systematic approach of human fact-checkers.
Stage 6: Justify the Verdict
Human readability and explainability are central in this final stage. Given the entire fact-checking report, DEFAME tasks the MLLM with generating a concise summary, highlighting key findings and critical evidence—including hyperlinks and images. The final output transparently documents the entire process, aiding verification and further human fact-checking.
4 Experiments
4.1 Datasets
We evaluate DEFAME on three well-known fact-checking datasets, representative of three distinct areas in fact-checking literature, and one novel benchmark comprised of claims after October 2023.
AVeriTeC (Schlichtkrull et al., 2024b) is a popular text-only real-world-based benchmark. The development set consists of $500$ claims: $305$ [Uncaptioned image]Refuted, $122$ [Uncaptioned image]Supported, $35$ [Uncaptioned image]NEI (Not Enough Information), and $38$ claims with the [Uncaptioned image]C/CP (Conflicting/Cherrypicking) label that designates claims with conflicting evidence or claims that are technically true but taken out of context. We retrieve evidence from the benchmark-complementary Knowledge Base (KB), which contains the necessary evidence along with approximately $1,000$ unrelated resources to simulate open web search. Thus, for AVeriTeC, the Web Search Tool does not utilize the Serper API but searches for documents via a semantic search. Each query to the KB is encoded using gte-base-en-v1.5 (Alibaba-NLP, 2024); the closest documents to the search query are retrieved via $k$-nearest neighbor. We evaluate DEFAME’s performance by computing accuracy over all $4$ classes.
Dataset GPT-4o GPT-4o mini LLaVA-1V
VERITE 84.5 67.1 59.3
MOCHEG 59.5 55.5 42.1
AVeriTeC 74.0 59.2 40.0
ClaimReview2024+ 67.0 53.7 36.9
Table 2: DEFAME accuracy for three different backbone models (GPT-4o, GPT-4o mini, LLaVA-1V) on the four benchmarks.
MOCHEG (Yao et al., 2023) is a multimodal dataset. Its textual claims are paired with multimodal text-image evidence and sourced from PolitiFact [^7] https://www.politifact.com/ and Snopes [^8] https://www.snopes.com/. The test set contains $2,001$ unique claims, out of which $1,689$ contain a final ruling. The ruling is used to assess the quality of generated justifications. Thus, we choose this subset as our test set. It includes $667$ [Uncaptioned image]Refuted, $522$ [Uncaptioned image]NEI, and $500$ [Uncaptioned image]Supported claims. We evaluate our performance using the micro $F1$-score for comparability with prior work while also reporting various metrics to assess justification quality in Appendix I.
The VERITE (Papadopoulos et al., 2024b) benchmark is designed for image-text claim verification focusing on the Out-Of-Context (OOC) scenario. It features claims with textual and image components, with some samples originating from fact-checking platforms and others obtained by swapping either an image or parts of a caption. We use the entire VERITE dataset totaling $1,001$ samples for evaluation.99913131313 claims were excluded due to issues in VERITE’s construction process, reducing the dataset from $1014$ to $1001$ samples. The samples are categorized into $338$ true, $325$ OOC, and $338$ miscaptioned claims [^10] These differ in the way they were constructed, but both result in out-of-context images.. We follow Papadopoulos et al. (2024a) and report “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned” and “[Uncaptioned image]True vs. [Uncaptioned image]OOC” accuracy. We further merge the two fabricated classes into one and report the “[Uncaptioned image]True vs. [Uncaptioned image]False” accuracy.
Finally, we construct ClaimReview2024+, a novel dataset with claims sourced from ClaimReview [^11] https://www.claimreviewproject.com/. This dataset aims to avoid the data leakage issue, i.e., MLLMs are likely pretrained on significant portions of claims and correlated fact-check articles from the previous three benchmarks. ClaimReview2024+ dataset consists of $300$ claims sourced after the October 2023 knowledge cutoff of GPT-4o. That way, we simulate a more realistic scenario where DEFAME is confronted with new events not encoded in the MLLM’s parameters. The dataset is subdivided into $160$ unimodal and $140$ multimodal claims, adhering to a four-class labeling scheme inspired by AVeriTeC: [Uncaptioned image]Supported, [Uncaptioned image]Refuted, [Uncaptioned image]Misleading, and [Uncaptioned image]NEI. This labeling also closely follows the classification by the fact-checking organizations that contribute to ClaimReview. Additional details about the dataset are included in Appendix J.
4.2 Model and Configuration
We chose GPT-4o and GPT-4o mini as the backbone of DEFAME since they are the current state-of-the-art MLLMs. To account for open-source MLLMs, we also test LLaVa-OneVision (1V) (Li et al., 2024a). DEFAME includes the MLLM without any fine-tuning, with temperature set to $0.01$ and top-$p$ to $0.9$ to control response diversity. We limit the number of images per scraped web page to a maximum of $32$ to avoid an excessive flood of images. DEFAME processes interleaved text-image inputs, preserving the original position of images within the text context, but any input exceeding the MLLM’s maximum context window of $128\,k$ tokens is truncated. Table 2 shows the performance of the different backbones.
4.3 Comparison to the State-of-the-Art
In Table 3, we present our evaluation results compared to State-of-the-Art (SOTA) methods and two baselines: GPT-4o, which directly generates a verdict (“Determine the claim’s veracity by picking one of the following decision options…”), and GPT-4o Chain-of-Thought (CoT), which also relies solely on parametric knowledge but is instructed to perform step-by-step reasoning before generating the verdict. DEFAME achieves an accuracy of $74.0\%$ on the AVeriTeC benchmark and surpasses the previous SOTA (Cao et al., 2023) that deploys a CoT LLM approach. With an overall accuracy of $84.5\%$ on VERITE, DEFAME ranks $32.7$ percentage points above prior best result (Papadopoulos et al., 2024a) in terms of “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned” accuracy. It performs competitively on the “[Uncaptioned image]True vs. [Uncaptioned image]OOC” accuracy, leading to an overall accuracy gain of $26.5$ percentage points. On MOCHEG, our framework achieves a micro $F1$-score of $59.5\%$, replacing the MetaSumPerceiver (Chen et al., 2024) as the new SOTA fact-checking system. Importantly, no prior work has demonstrated the ability to simultaneously address such diverse tasks as we do here.
Model AVeriTeC MOCHEG VERITE CR+
Acc F1 T/OOC T/MC T/F Acc
CFR (2024) 60.0 - - - - -
GPT-CoT (2023) 65.6 - - - - -
LVLM4FV (2024) - 45.1 - - - -
MetaSum (2024) - 48.6 - - - -
CHASMA (2024b) - - 74.4* 59.3* 52.1 -
AITR (2024a) - - 82.7 51.8 58.0* -
GPT-4o 61.4 53.5 68.9 71.6 78.0 34.7
GPT-4o CoT 60.6 49.6 73.6 76.2 79.7 36.6
DEFAME 74.0 59.5 79.8 84.5 84.5 67.0
Table 3: Comparison of our method with prior works across datasets. Best scores are in bold, second best are underlined. Values marked with * had to be derived from the reported numbers. CR+ = ClaimReview2024+.
The results on our new ClaimReview2024+ dataset challenge the notion that vanilla LLMs are reliable fact-checking systems. While performance on established benchmarks is strong, ClaimReview2024+ reveals a drastic drop for the two GPT-4o baselines, whereas DEFAME maintains accuracy comparable to its performance on AVeriTeC. This suggests that DEFAME’s evidence retrieval mitigates the temporal dependence of its backbone model. The confusion matrices (Figure 4) illustrate distinct error patterns across models on the ClaimReview2024+ dataset. Pure GPT-4o overpredicts [Uncaptioned image]NEI but outperforms its Chain-of-Thought counterpart in identifying [Uncaptioned image]Refuted claims. GPT-4o CoT also acknowledges the lack of evidence but shows a stronger tendency to commit to [Uncaptioned image]Misleading. In contrast, DEFAME exhibits a more balanced confusion matrix, with its primary challenge lying in differentiating [Uncaptioned image]Misleading and [Uncaptioned image]Refuted —a relatively uncritical ambiguity among the possible misclassifications. Its result on the [Uncaptioned image]Supported class is far superior to the two baselines.
Figure 3: Example claim from ClaimReview2024+ with verdict and justification by GPT-4o CoT and DEFAME.
For example, when confronted with the claim, “Slovakian Prime Minister Robert Fico being dragged into a car after being shot,” (Fig. 3) the GPT-4o CoT baseline dismisses it as false due to a lack of evidence. In contrast, DEFAME conducts a reverse image search and web search, finds the image, including the corresponding articles on CNN, Vatican News, and Al Jazeera, and concludes that “the evidence confirms that Fico was indeed shot on May 15, 2024, in Handlová, Slovakia, and subsequently moved into a car after the incident.” Even though the geolocation suggested Poland and the Czech Republic, DEFAME correctly marginalized this finding and predicted the correct verdict.
Further analysis is provided in Appendix D.
(a) ClaimReview - GPT-4o
(b) ClaimReview - GPT-4o CoT
(c) ClaimReview - DEFAME
Figure 4: Confusion matrices for GPT-4o, GPT-4o CoT, and DEFAME on the ClaimReview2024+ dataset.
4.4 Ablation Study
MOCHEG VERITE
Model Variant F1 T/F (Acc.)
DEFAME 59.5 84.5
w/o Geolocation 58.3 80.6
w/o Reverse Search 58.2 73.7
w/o Image Search 57.8 81.4
w/o Web Search 42.0 81.8
Single Turn 47.7 82.8
Static Actions 58.7 83.0
w/o Develop 57.4 83.8
Unimodal Develop 56.1 82.0
Table 4: Ablation study results for different model variants on the MOCHEG and VERITE datasets. “True vs. False(= MC + OOC)” accuracy is reported for VERITE. Best scores are marked in bold, second best are underlined.
We conduct an ablation study with reduced DEFAME variants to investigate the contributions of various components and capabilities of DEFAME to its overall performance:
Tool Ablations: We assess the contribution of each of the four tools by individually removing them from the tool pool.
Single Turn: This version is restricted to pass all stages only once, i.e., no ability to delve deeper into findings.
Static Actions: This variant fixes all actions into a static action schedule. Each tool is executed exactly once, bypassing the dynamic planning stage.
W/o Develop Stage: Is intermediate reasoning useful? We test a variant that excludes the Develop Stage, jumping from the evidence retrieval immediately to the judgment.
Unimodal Develop Stage: Does it help to carry images through the stages and iterations, or is textual reasoning sufficient? This part-unimodal variant can retrieve evidence images but cannot pass them on to future stages.
The results of the ablation study are displayed in Table 4. Analyzing the contribution of each tool, we find that DEFAME benefits from all four tools. While Web Search and Image Search are important for both datasets, Geolocation and Reverse Image Search yield little gain on MOCHEG. This is likely caused by the benchmark’s focus on textual claims. On the other hand, the images included in the VERITE claims can be immediately scrutinized with Geolocation and Reverse Image Search tools, thereby increasing these tools’ relevance for text-image claims.
The single-turn variant of DEFAME lags behind the multi-turn version, especially on the MOCHEG benchmark. This suggests that for MOCHEG follow-up retrievals are important. Our findings also highlight the significant impact of the Develop Stage and the integration of visual evidence from tool outputs into the fact-check report. This multimodal enhancement yields a measurable performance boost, improving results by around 3 percentage points on MOCHEG and VERITE. Additionally, our experiments indicate that a static action routine leads to inferior accuracy as well as significantly higher computational costs (cf. Appendix B).
4.5 Adaptation to the AVeriTeC Challenge
Figure 5: Top 8 systems on the AVeriTeC Challenge test set, ranked by AVeriTeC score (in %).
The AVeriTeC dataset has recently run a challenge, where methods compete in terms of a new metric, AVeriTeC Score, aimed to capture accuracy and evidence correctness, see Schlichtkrull et al. (2024b) for details. To compute this metric, we adapt our approach to include question-answer (QA) pairs since the score measures the similarity between generated and gold QA pairs. See Appendix C for more details. Our adapted approach (coined DEFAME-QA) achieves stellar results in the AVeriTeC Challenge, shown in Table 5). Notably, both the original DEFAME and the adapted DEFAME-QA achieve the same accuracy on the development set.
4.6 Explainability Quality and Human Evaluation
To further measure the quality of gathered evidence and to assess our framework’s vulnerability to hallucinations, we conducted a human evaluation, focusing two aspects:
Coherence: The fact-check maintains a logical and meaningful flow. There are no contradictions or gaps that disrupt the overall coherence of the report.
Completeness: The verdict is sufficiently justified. That is, the included evidence allows one to derive the verdict.
Figure 6: Human assessment of Coherence and Completeness of DEFAME’s vs. GPT + CoT’s fact-checking reports
Participants were asked to rate the fact-checking reports w.r.t. the above criteria on a Likert scale from $1$ to $5$. In total, we collect $185$ ratings over $36$ claims sampled randomly but balanced from VERITE, MOCHEG, and AVeriTeC. Please refer to Appendix H for further details on the experiment. The results are shown in Fig. 6. The CoT baseline and DEFAME show no significant difference in output coherence, i.e., both produce mostly logical reports. This is expected, considering the current state of zero-shot LLMs in text generation. However, the two differ significantly in their ability to justify their verdicts. The results imply that DEFAME provides better justifications compared to bare MLLM prompting. This finding further challenges recent studies (Geng et al., 2024; Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e) that suggest MLLMs can perform fact-checking without retrieving external evidence.
4.7 Failure Analysis
We analyzed 20 VERITE cases where DEFAME mislabeled a sample and identified three common failure modes. (1) Premature Judgment, the most frequent one, occurs when DEFAME commits to a verdict before gathering sufficient evidence. Often, it predicts [Uncaptioned image]Supported as soon as it finds any image matching the claim text, even if it is not the one in the claim. (2) Failed Evidence Retrieval accounts for 25% of errors, where DEFAME exhausts all three retrieval attempts but fails to find enough evidence. This happens mainly when evidence is inaccessible (e.g., when API access is required) or embedded in a video. (3) Wrong Ground Truth: In three cases, our inspection revealed annotation errors, meaning DEFAME’s prediction was actually correct. In one instance, DEFAME “over-summarized” the evidence, omitting critical details.
5 Discussion
While DEFAME works with any MLLM backbone, the backbone’s capabilities directly influence the quality of the resulting fact-check. As MLLMs continue to evolve, DEFAME’s performance is expected to improve correspondingly. Moreover, its modular, in-context-learning based design enables DEFAME to incorporate further tools beyond those demonstrated in this work. Still, DEFAME has several limitations.
Credibility of External Evidence: Our reliance on search engines introduces the risk of incorporating unreliable information, as search results may include leaked or biased content (Chrysidis et al., 2024). While some approaches assess source credibility using third-party ratings  (Zhou et al., 2024), our framework relies on diversification and the search engine’s internal trustworthiness checks121212blog.google/products/search/how-google-delivers-reliable-information-search/. Introducing external ratings could strengthen verification rigor.
System Stability and Sensitivity: Our web scraping process is prone to instability due to restricted access and large document size. Moreover, open-source models like Llava exhibit formatting sensitivity, where minor prompt variations impact response quality. Ensuring robust scraping and prompt formatting stability is crucial for reliable outputs.
Hallucination: We acknowledge that hallucinations are a risk in DEFAME as they are inherent to LLMs. However, both qualitative analysis and human evaluation do not indicate substantial amounts of hallucinations. Still, we agree that the role of hallucination in DEFAME must be analyzed more closely in future work.
6 Conclusion
We presented DEFAME, a comprehensive zero-shot framework for multimodal fact-checking that integrates MLLMs with external tools to address the limitations of traditional automated fact-checking approaches. Our framework explainably grounds its analysis in verifiable data by combining MLLM-driven reasoning with external multimodal evidence sources, geolocation, reverse image search, and dynamic web searches. We demonstrated that DEFAME achieves new state-of-the-art results across multiple diverse benchmarks, including an own leakage-free benchmark, highlighting its capability to navigate unseen real-world claims. While limitations remain, our system represents a significant advancement in the field, providing a versatile foundation for future developments in automated multimodal fact-checking.
7 Impact Statement
In order to scalably debunk the masses of mis- and disinformation—which indisputably pose a strong threat to social cohesion—automating fact-checking is inevitable. Although the results do not sufficiently warrant real-world application yet, DEFAME’s introduction marks a pivotal step in addressing that challenge.
DEFAME contributes shifting the paradigm toward explainable and verifiable AI systems. Its ability to provide structured, evidence-backed justifications ensures that AFC does not become a black-box authority but rather an augmentation tool that strengthens public resilience against misinformation. By embedding transparency and iterative verification into its core process, DEFAME helps counter the skepticism toward AI-driven moderation while simultaneously reducing the burden on human fact-checkers.
However, the widespread deployment of such systems also raises ethical and societal concerns. Recently, in public discussion, professional fact-checkers have widely faced accusations of being biased and, therefore, unreliable. Automated fact-checking, even when well-calibrated, can also shape narratives by prioritizing certain sources or by amplifying institutional biases inherent in retrieval mechanisms or parametric knowledge. A recent example for governmental censorship in LLMs is the DeepSeek LLM family. Current research efforts strive to de-bias LLMs, which DEFAME will automatically benefit from thanks to its backbone-agnostic design. Within the strongly polarized public discussions, AFC systems may emerge as a bipartisan information source. Their reliability inherently depends on the quality of the backbone, which is difficult to rigorously analyze.
There is also a risk that governments or platforms might misuse systems like DEFAME for overreach in content moderation, potentially stifling dissent or critical journalism under the guise of misinformation control. To mitigate this, it is crucial to ensure accountability, source diversity, and contestability—allowing users to challenge and scrutinize automated decisions. Additionally, while DEFAME may improve trust in digital information, its success will ultimately depend on how it is integrated into broader fact-checking ecosystems and content moderation, where human oversight remains essential. The current debate strikingly shows that fact-checking (especially of social media content) sits on the verge between public safety and perceived censorship of free speech. Therefore, the real impact of DEFAME lies not just in its technical contributions but in how humans will ultimately use it.
8 Acknowledgments
The research was partially funded by a LOEWE-Start-Professur (LOEWE/4b//519/05.01.002-(0006)/94), LOEWE-Spitzen-Professur (LOEWE/ 4a//519/05.00.002-(0010)/93) and an Alexander von Humboldt Professorship in Multimodal Reliable AI, sponsored by Germany’s Federal Ministry for Education and Research.
Appendix
The appendix provides an in-depth extension of the main paper, showcasing additional examples, analyses, and details that complement the findings discussed. Section A outlines excluded domains to ensure a fair and realistic evaluation setting. We include information on compute and cost in Section B. Details about our adaptations for the AVeriTeC challenge are given in Section C. A short analysis of the confusions committed by the CoT variant and DEFAME is shown in Section D. Section E includes the prompts used in DEFAME. Section F identifies wrongly annotated samples in the VERITE dataset. Section G discusses cases with outdated ground truth in the MOCHEG dataset and their implications. Section H specifies our human evaluation of the fact-checking report. In Section  I, we analyze the justifications generated by DEFAME and scrutinize current evaluation metrics. We present our ClaimReview2024+ dataset in Section J. Section K illustrates an example fact-checking report produced by DEFAME. Lastly, Section L delves into failure cases, highlighting common pitfalls and their underlying causes.
Appendix A Excluded Domains
To maintain a realistic and fair setting, we exclude all major fact-checking organizations we know from all web search results. Table 5 shows the corresponding list of domains. Additionally, several platforms forbid (direct) automatic access to their web pages, cf. Table 6. Any URL with a sub-string matching the domains and URLs in the Tables mentioned earlier is removed from the search results and, hence, ignored by the fact-check.
Excluded Fact-Checking URLs
snopes.com
politifact.com
factcheck.org
truthorfiction.com
fullfact.org
leadstories.com
hoax-slayer.net
checkyourfact.com
reuters.com/fact-check
reuters.com/article/fact-check
apnews.com/APFactCheck
factcheck.afp.com
poynter.org
factcheck.ge
vishvasnews.com
boomlive.in
altnews.in
thequint.com/news/webqoof
factcheck.kz
Table 5: List of excluded URLs to maintain a fair and realistic fact-checking setting.
Unsupported Domains
facebook.com
twitter.com
x.com
instagram.com
youtube.com
tiktok.com
reddit.com
ebay.com
microsoft.com
researchhub.com
pinterest.com
irs.gov
Table 6: List of excluded domains due to bot traffic restrictions.
Model Variant
Time min / Claim
LLM API Cost ¢ / Claim
Search API Cost ¢ / Claim
# Search Queries Searches / Claim
# Input Tokens Processed Tokens / Claim
# Output Tokens Generated Tokens / Claim
DEFAME:
- GPT-4o 4:14 13 $<$1 2.2 48K 921
- GPT-4o mini 3:13 $<$1 $<$1 3.1 54K 1300
- LLaVA-1V 1:23 - $<$1 0.9 14K 794
w/o Geolocation 3:12 11 $<$1 2.3 40K 910
w/o Reverse Search 2:37 8 $<$1 2.3 28K 828
w/o Image Search 2:57 11 $<$1 2.2 40K 909
w/o Web Search 2:30 10 $<$1 2.2 37K 761
Single Turn 2:47 10 $<$1 2.0 38K 858
Static Actions 5:11 14 $<$1 5.2 70K 1100
w/o Develop 1:56 10 $<$1 2.1 42K 710
Unimodal Develop 3:15 10 $<$1 2.3 41K 865
Pure GPT-4o 0:06 $<$1 0 0.0 2.2K 67
Chain-of-Thought 0:09 $<$1 0 0.0 2.3K 177
Table 7: Usage and computational cost statistics of DEFAME and all considered variants on VERITE. Time values are rough estimates and object to high variance due to hardware usage by other processes.
Appendix B Details on Compute and Cost
As the GPT models are available only via OpenAI’s API, most of our computation happens externally. On our end, we employed four NVIDIA A100-80GB GPUs in order to execute LLaVA-1V and GeoCLIP. All other processing was performed on 32 AMD EPYC 7313 16-core CPUs. Since fact-checks on VERITE claims are multimodal from the start, we chose VERITE as a representative surrogate for the setting of multimodal AFC. We thus report usage statistics on VERITE in Table 7.
As expected, the naive baselines (Pure GPT-4o and CoT) require the least amount of resources. But, as discussed in Section 4, this is due to the absence of any external evidence and comes at the cost of lower accuracy. DEFAME with GPT-4o mini and LLaVA-1V as the backbone MLLM runs faster and cheaper than DEFAME with GPT-4o, but yields decreased accuracy as well. Surprisingly, LLaVA-1V fact-checks faster than the GPT models. We found the reason to lie in LLaVA-1V’s inability to correctly format the proposed actions, yielding a smaller number of executed actions and, thus, shorter fact-checks overall.
Compared to human fact-checking experts—who invest about an entire working day to debunk a claim and write a corresponding article (Hassan et al., 2015)DEFAME’s fee of \$$\$0.13$ per claim is cheap. However, DEFAME’s output quality does not match human fact-checkers yet. Thus, DEFAME could serve as a cost-effective assistant to aid human fact-checkers. Theoretically, social media platforms could also use DEFAME to cheaply mass-check larger amounts of claims posted online. However, depending on the number of claims—which, on social media, arguably exceed millions per day—DEFAME could become expensive. Claim filtering approaches would be needed to narrow down the claims to the most check-worthy ones in order scale DEFAME up to larger amounts.
Appendix C Adaptation to the AVeriTeC Challenge
The AVeriTeC challenge evaluates fact-checking quality using the AVeriTeC Score that compares model-generated question-answer (QA) pairs with gold QA pairs provided in the benchmark (Schlichtkrull et al., 2024b). To perform well, a method must effectively identify and address the same (or very similar) questions and answers posed by the benchmark annotators.
To align with this evaluation, we introduce an extension called DEFAME-QA. This adaptation begins by generating 10 key questions designed to probe the claim’s veracity. The Planner then proposes targeted search queries for each question and applies the Web-Search tool to retrieve up to 5 relevant search results. Using the retrieved evidence, the LLM backbone attempts to answer the questions systematically. Finally, the system synthesizes the answers into a coherent verdict, ensuring the reasoning is grounded in the collected evidence. The resulting QA pairs are evaluated using the AVeriTeC Score, showcasing DEFAME-QA’s alignment with the benchmark’s evaluation criteria while maintaining its structured, evidence-driven methodology.
Appendix D Model Confusions
Figure 7 shows the confusion matrices for all three backbones on the four benchmarks, respectively.
A closer look at Figures 7(a), 7(b) and 7(c) reveals that GPT-4o with and without Chain-of-Thought overpredicts [Uncaptioned image]OOC while DEFAME’s confusions are more balanced. Surprisingly, incorporating Chain-of-Thought prompting hurts the performance on MOCHEG (see Figures 7(d), 7(e) and Table  3). According to the confusion matrices the introduction of Chain-of-Thought makes the model more unsure, leaning towards [Uncaptioned image]NEI compared to the Pure GPT-4o. In contrast, DEFAME predicts too confidently even when the ground truth implies insufficient information. A qualitative analysis of the failure cases reveals that, in several cases, the ground truth explanation is no longer up-to-date (see Sections 4.7 and Appendix F).
On AVeriTeC (Figures 7(g), 7(h), and 7(i)), we observe the opposite behavior with the GPT-4o variants overpredicting [Uncaptioned image]NEI compared to our framework. Lastly, DEFAME’s false predictions on the ClaimReview2024+ dataset appear balanced in Figure 4 while the baselines lack the necessary evidence to make correct veracity predictions, very often defaulting to [Uncaptioned image]NEI.
(a) VERITE - Pure GPT-4o
(b) VERITE - Chain-of-Thought
(c) VERITE - DEFAME
(d) MOCHEG - Pure GPT-4o
(e) MOCHEG - Chain-of-Thought
(f) MOCHEG - DEFAME
(g) AVeriTeC - Pure GPT-4o
(h) AVeriTeC - Chain-of-Thought
(i) AVeriTeC - DEFAME
Figure 7: Confusion matrices for Pure GPT-4o, Chain-of-Thought, and DEFAME across the VERITE, MOCHEG, and AVeriTeC datasets.
Appendix E LLM Prompts
Each stage of the DEFAME framework is guided by a tailored prompt. These prompts are constructed from prompt templates, where values enclosed in [] serve as placeholders. During execution, these placeholders are dynamically replaced with the corresponding variables. This process ensures that the prompt is specific to the current task and context, as illustrated in E.1 and E.2.
In subsequent sections, we present the templates for the remaining four stages of the DEFAME framework. Each template includes detailed explanations of its placeholders. A key placeholder is [Record], which is present in every prompt. This generic placeholder provides the current state of the fact-checking report, consolidating the claim, evidence, and findings gathered so far, ensuring that the LLM operates within the relevant context.
Some prompts require specific values to be returned by the LLM, such as the verdict in the Judge Prompt or the proposed actions in the Plan Prompt. In these cases, both the expected value and its format are explicitly defined within the prompt to guide the LLM’s response. Value-specific fallback mechanisms are employed to enhance robustness. These mechanisms, primarily based on regular expressions tailored to observed failure modes, ensure that the required values can still be reliably extracted, even if the LLM deviates from the expected format or introduces minor inconsistencies.
For multimodal LLM inputs and outputs, images are integrated into prompts through a referencing system. When an image reference (e.g., image:k) is encountered, the system inserts a corresponding image block, including the Base64-encoded representation of the image. The LLM references these images using the same identifier, maintaining a consistent link between visual and textual elements.
The prompt templates in DEFAME are dataset-agnostic, enabling use across benchmarks with minimal adaptations. Dataset-specific changes are limited to [Extra Rules] in the Plan and Judge Prompts. MOCHEG requires no additional rules, while AVeriTeC includes a guideline to avoid the ”argument from ignorance” fallacy, ensuring unsupported claims are labeled as Not Enough Information. In VERITE, detailed instructions are needed for the [Uncaptioned image]OOC class, which includes samples generated in two ways: true images with altered captions (formerly [Uncaptioned image]Miscaptioned) and true captions with unrelated images. These rules address dataset-specific nuances while maintaining a consistent framework.
E.1 Plan Prompt Template
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as many as needed. Do not propose similar or previously used actions.
[Extra Rules]
[Valid Actions]
[Examples]
[Record]
Your Actions:
  • [Extra Rules]: Contains benchmark-specific planning guidelines that the Planner must follow when selecting actions. These rules are tailored to the requirements of individual datasets or evaluation scenarios, ensuring that the Planner adheres to task-specific constraints or priorities.
  • [Valid Actions]: Represents the set of actions available to the Planner at a given stage. The list of valid actions is dynamically adapted to avoid reusing the same action unnecessarily.
  • [Examples]: Provides in-context examples that demonstrate how to use actions with the correct format. These examples illustrate the structure and logic behind action proposals to guide the Planner.
E.2 Finalized Plan Prompt
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as much as needed. Do not propose similar or previously used actions. Consider Both Modalities Equally: Avoid focusing too much on one modality at the expense of the other, but always check whether the text claim is true or false. Compare Image and Caption: Verify the context of the image and caption.
Valid Actions: geolocate: Determine the country where an image was taken by providing an image ID. reverse_search: Perform a reverse image search on the web for similar images. web_search: Run an open web search for related webpages. image_search: Retrieve related images for a given query.
Examples: geolocate(<image:k>) reverse_search(<image:k>) web_search("New Zealand Food Bill 2020") image_search("China officials white suits carry people")
Record: Claim: “<image:1232>Image of a bus powered by compressed natural gas, bursting into flames in Italy.”
Your Actions:
E.3 Summarize Prompt Template
Instructions In order to find evidence that helps your fact-check, you just ran a web search, which yielded a Search Result. Your task right now is to summarize the Search Result. What to include: Information that might be useful for the fact-check (see Record). Relevant images (refer to images by inserting their reference in the format <image:k>). If available: the release date as well as the author or the publisher (e.g., the media company) of the search result. Do NOT include: Advertisements. Any other information unrelated to the Record or the Claim. Additional Rules: Do not add any additional information besides the information in the Search Result. If the Search Result doesn’t contain any relevant information for the fact-checking work, simply print one word in capital letters: NONE. Keep your writing style consistent with the provided Examples. Try to filter out relevant information even if the Search Result is in a different language.
[Examples]
[Record]
[Search_Result]
Your Summary:
  • [Examples]: Provides 3 in-context examples that demonstrate how to write concise summaries, incorporating relevant images, key insights, and links to sources using Markdown notation. One of the examples shows a case where the search result is irrelevant, guiding the model to return NONE instead of a summary when no useful information is found.
  • [Search_Result]: Refers to the search result retrieved by a Search tool. This includes the text content scraped from the webpage using the Firecrawl web scraper, the title of the page, any hyperlinks found within the content, images included on the page, and the source URL.
E.4 Develop Prompt Template
Instructions You just retrieved new Evidence. Now, analyze the Claim’s veracity using the evidence. Always adhere to the following rules: Focus on developing new insights. Do not repeat larger parts from the Record. Do not restate the Claim. Write down your thoughts step-by-step. Whenever necessary, you may elaborate in more detail. Depending on the topic’s complexity, invest one to three paragraphs. The fewer, the better. If you find that there is insufficient information to verify the Claim, explicitly state what information is missing. If you cite web sources, always refer to them by including their URL as a Markdown hyperlink. Use information only from the recorded evidence: Avoid inserting information that is not implied by the evidence. You may use commonsense knowledge, though.
[Record]
Your Analysis:
E.5 Judge Prompt Template
Instructions Determine the Claim’s veracity by following these steps: 1. Briefly summarize the key insights from the fact-check (see Record) in at most one paragraph. 2. Write one paragraph about which one of the Decision Options applies best. Include the most appropriate decision option at the end and enclose it in backticks like ‘this‘.
[Extra Rules]
[Decision Options]
[Record]
Your Judgement:
  • [Extra Rules]: Contains benchmark-specific rules or additional constraints that guide the judgment process. These rules are designed to increase the model’s understanding of the different classes present in the corresponding benchmark, ensuring consistent verdicts.
  • [Decision Options]: Lists the possible labels or verdicts that can be assigned to the claim, along with a short description of each label. These descriptions provide additional context to help the model accurately differentiate between the available options.
E.6 Justify Prompt Template
Instructions You are provided with the record of a fact-check. It contains the Claim to be verified and documentation of all the fact-checking work along with the gathered evidence. Your task is to summarize the fact-check. That is, you provide a concise, one-paragraph justification for the final VERDICT based on the knowledge from the Record. Note: Be truthful, brief, and do not add any additional information besides the information given in the Record. Link key sources in your summary. Use Markdown notation for that. You may link them in-line. Don’t state the Claim again. Rather focus on the key insights of the fact-check. Simply print just the summary.
[Record]
Summary:
Appendix F Wrongly Annotated VERITE Samples
During the qualitative analysis of $20$ mispredicted VERITE instances, we encountered $3$ cases we argue to be wrongly annotated. Figure 8 shows the corresponding claims and (wrong) annotations. The VERITE annotation classifies the first claim ($753$) as [Uncaptioned image]Supported. However, DEFAME found the image to be captured not in September but in July 2018, citing a fact-check by USA Today [^13] https://web.archive.org/web/20220526225427/https://eu.usatoday.com/story/news/factcheck/2022/05/26/fact-check-photo-ginni-thomas-expensive-wine-2018/9910097002/ from May 2022. Manually investigating this case further, we find that USA Today refers to an article on Mediaite [^14] https://web.archive.org/web/20180906222511/https://www.mediaite.com/online/exclusive-clarence-thomas-wife-hired-ex-tpusa-staffer-known-for-saying-i-hate-blacks/ from Sept 6th, 2018, indeed stating that “in July, Clanton shared a photo on Instagram of herself, Thomas, and Ginni Thomas having a ‘great weekend’ together,” showing the screenshot of the respective Instagram post. Hence, the actual correct label for this claim should be [Uncaptioned image]OOC.
The wrong annotation can be seen even more clearly for the claims $249$ and $250$. DEFAME’s reverse search yielded multiple credible sources, including an article by CBS News1515footnotemark: 15, consistently reporting about a gathering of $75$ people at the “Area 51” military airbase in Sept 2019. 161616www.cbsnews.com/news/storm-area-51-hoax-draws-hundreds-events-outside-secretive-us-base-today-2019-09-20-live-updates/ The sources use the claim’s photo along with other similar images showing the apparently same event. According to the VERITE annotation, the photo shows a completely different event, contradicting the evidence. Consequently, the provided labels for both claims are clearly “switched.” Since we analyzed only $20$ samples, there are likely more such wrongly annotated samples, penalizing DEFAME’s accuracy where it should not be. Hence, the actual accuracy of DEFAME is slightly higher than measured.
Claim # 753 “An image of Supreme Court Justice Clarence Thomas with his wife, Ginni Thomas, holding a bottle of wine was captured in September 2018 after hiring Crystal Clanton to assist her media ventures.” Refer to caption Annotation Refer to captionSupported Actually Refer to captionOOC
Claim # 250 “Image shows a crowd of people at the ’Area 51 Raid’ in September 2019.” Refer to caption Annotation Refer to captionOOC Actually Refer to captionSupported
Claim # 249 “Image shows a crowd of people during a ’crusade’ by the religious group ’In His Name Ministries’ in November 2014 in Nikomazi, South Africa.” Refer to caption Annotation Refer to captionSupported Actually Refer to captionOOC
Figure 8: Three faulty VERITE instances identified during qualitative analysis.
Appendix G Outdated Ground Truths and Temporal Dependence in MOCHEG
A qualitative analysis of failure cases in the MOCHEG dataset reveals that some ground truth explanations are no longer accurate or up-to-date, potentially affecting model evaluations (see Section 4.7). This issue arises when real-world developments render previously valid ground truths obsolete.
For instance, consider the claim:
“A company is developing a lab-grown chicken nugget made from feathers.”
This was classified as Not Enough Information by Snopes, with the explanation that it is undetermined when these lab-grown nuggets will hit store shelves. [^17] https://www.snopes.com/fact-check/chicken-nuggets-feather-cells However, since the creation of the benchmark, lab-grown meat has become commercially available in parts of the world. [^18] https://www.theguardian.com/food/2020/dec/07/lab-grown-chicken-tastes-like-chicken-but-the-feeling-when-eating-it-is-more-complicated
Other examples of claims with temporal dependence include:
  • Claim: “Al Gore’s residence uses considerably more energy than the average American home.”
  • Claim: “Oreos have a new Lady Gaga-themed cookie.”
  • Claim: “Says Anthony Fauci will make millions off new book.”
Such claims rely on a specific temporal context that may no longer be accurate as time progresses. Including such claims without temporal markers risks introducing inaccuracies into evaluation. To address the challenge of temporal dependence in fact-checking benchmarks, we propose the following alternatives:
  1. 1.
    Include Timestamps: Ensure that datasets include clear timestamps for both the claim and the associated ground truth explanation, allowing systems to account for the temporal context.
  2. 2.
    Filter Out Time-Sensitive Claims: Exclude claims with high temporal sensitivity from the dataset to avoid potential inconsistencies over time.
  3. 3.
    Periodic Updates: Regularly update benchmarks to reflect evolving ground truths, ensuring their continued relevance.
  4. 4.
    Temporal Validity Check: Integrate a pre-processing step to verify whether the ground truth explanations remain consistent with current knowledge before evaluation.
Appendix H Details on the Human Evaluation
We ensure a minimum number of $5$ evaluations per claim and include random samples from the Chain-of-Thought baseline for comparison. (We post-process the baseline outputs to match the format of DEFAME’s output to “disguise” it.) In total, $154$ of the submissions assess the outputs from DEFAME while $31$ assess the baseline. An example of the report format is included in Appendix K. We assess the difference in Completeness scores between the DEFAME and CoT LLM groups using the Mann-Whitney U Test. This non-parametric test was chosen due to the non-normal distribution of completeness scores. The Mann-Whitney U Test yields a $p$-value of approximately $9.1\times 10^{-9}$, deeming the findings statistically significant. All Participants have a higher degree in education.
To provide further insights, we include a direct comparison of a fact-checking report generated by DEFAME and one from the Chain-of-Thought baseline for the same claim (see Figures  9 and  10). This example illustrates the key differences observed during the evaluation, particularly in the Completeness dimension. While both reports maintain coherence in structure and logical flow, the DEFAME report explicitly links its verdict to multiple pieces of evidence, providing a clear justification. In contrast, the CoT report relies heavily on parametric reasoning, lacking grounded evidence to support its conclusions.
While our human evaluation highlights DEFAME’s strengths, we acknowledge certain limitations, such as the relatively small number of claims evaluated.
Figure 9: Fact-checking report by DEFAME, presented in the Human Evaluation.
Figure 10: Fact-checking report by the Chain-of-Thought baseline, presented in the Human Evaluation.
Appendix I Automatic Justification Evaluation
We also compare the generated justifications on the MOCHEG dataset to the gold rulings in Table 8. Note that our competitors here are fine-tuned on the gold rulings (except the top row and CoT). The vanilla zero-shot performance of DEFAME trails behind the other methods. We experimented by providing a single in-context example to convey the expected style of the rulings. It leads to a significant improvement in all metrics, even compared to the state-of-the-art fine-tuned models. Nonetheless, rather than speaking for much improved explanatory power, this shows the shortcomings of the automatic metrics, e.g., their sensitivity to specific n-grams and wording.
Model ROUGE L BLEU BERTScore
Best w/o FT (Yao et al., 2023) 17.18 7.38 83.95
FT (Chen et al., 2024) 24.60 11.40 88.10
FT (Yao et al., 2023) 24.83 10.08 86.95
DEFAME 18.72 3.20 85.89
DEFAME 1-shot 25.37 7.31 87.42
Table 8: Performance Comparison of Explanation Generation (in %). Best scores are marked in bold, second best are underlined.
Appendix J ClaimReview2024+
To construct the ClaimReview2024+ dataset, we scraped 300 claims from the ClaimReview API provided by Google [^19] https://toolbox.google.com/factcheck/apis, covering the period from November 1, 2024, to January 18, 2024. The dataset consists of 129 [Uncaptioned image]Refuted, 89 [Uncaptioned image]Supported, 61 [Uncaptioned image]Misleading, and 21 [Uncaptioned image]NEI claims, with 160 being text-only and 140 paired with a corresponding image. While the labels follow the annotations provided by fact-checking websites, adjustments were made to align them with our taxonomy.
We applied a combination of automated and manual steps to standardize labeling. Automated classification was based on patterns in the original annotations, while ambiguous or conflicting cases were manually reviewed for consistency. Additionally, some claims were cleaned to remove explicit indications of veracity in the text. For example, a claim like ”Fact Check: Image from Pakistan falsely claims to show Punjab police using drugs” was reworded to ”Image from Pakistan shows Punjab police using drugs” to avoid embedding the claim’s outcome in its phrasing.
Finally, while this dataset was designed to contain claims novel to GPT-4o, we acknowledge the possibility that some claims may have appeared in earlier forms and were later re-evaluated by fact-checking websites before being picked up by ClaimReview.
Appendix K Example Fact-Checking Report
Figures 11 and 12 display a 2-page fact-checking report as returned by DEFAME, including a correct veracity prediction. We rendered the Markdown output into a PDF, including the referenced images. This and the following reports also include hyperlinks [^20] For technical reasons, the hyperlinks are not preserved in this PDF., referencing the used resources.
Figure 11: Exemplary fact-check report with a correct prediction, page 1 of 2.
Figure 12: Exemplary fact-check report with a correct prediction, page 2 of 2.
Appendix L Examples of Failure Cases
The qualitative analysis of mispredicted VERITE instances uncovered two common failure modes attributed to DEFAME: premature judgment and failed evidence retrieval. See Figure 13 for an example of a premature judgment. While it is true that there exists a photo showing the “Air Force Officers’ Spouses’ Club” taken on April 24, 2018212121https://web.archive.org/web/20241116170848/https://taskandpurpose.com/military-life/viral-military-spouse-tweet-photo/, it is not the photo depicted in the claim. DEFAME missed comparing the claim photo with evidence photos before judging the veracity.
Additionally, Figures 14 and 15 show a case where the retrieval of evidence (from either (reverse) image search or web search) was unsuccessful, resulting in a wrong prediction. Manual inspection reveals that according to Snopes [^22] https://www.snopes.com/fact-check/sharks-on-power-lines-hurricane/, the origin is a news video accessible on YouTube. However, both Snopes and YouTube are excluded from DEFAME’s search results. Apart from that, manually using all three web search tools yields only more excluded or unrelated results.
Figure 13: Report of a fact-check which ended in a wrong prediction due to a premature judgment.
Figure 14: Report of a fact-check (page 1 of 2), which ended in a wrong prediction due to the failed retrieval of evidence.
Figure 15: Report of a fact-check (page 2 of 2), which ended in a wrong prediction due to the failed retrieval of evidence.